# Quantum simulating continuum field theories with large-spin lattice models

### v1

@@ IV.2 Preparing a semi-classical soliton
To achieve the desired initial state, we first calculate the ground state of the sG Hamiltonian $\hat{H}_{\text{sG}}$ at finite $\beta^{2}$. For our microscopic implementation (at $\kappa=1$), this state corresponds to a correlated chain of spins pointing in $+x$-direction, as $\langle\beta\hat{\varphi}\rangle=\arg{(\langle\hat{J}_{+}\rangle/\sqrt{J(J+1)})}=0$. For a static soliton with $v=0$ and $\gamma=1$, it is now sufficient to imprint the desired phase profile by applying site-dependent $z$-rotations $\hat{U}_{\text{p}}^{(i)}=\exp{\left(-i\beta\varphi_{\text{s}}^{(i)}\hat{J}_{z}^{(i)}\right)}$, with $\varphi_{\text{s}}^{(i)}=\varphi_{\text{s}}(x=i\cdot a,t=0)$ the value of the phase profile at site $i$ [see Fig. 5(a)].
Figure 5: Soliton preparation and dynamics. (a)-(b) Preparation of the semi-classical soliton state: (a) we first apply site-dependent $z$-rotations with an angle $\beta\varphi_{\mathrm{s}}^{(i)}$ (green shaded area) to rotate in the equatorial plane the spins (initially pointing in $+x$-direction). (b) We then apply rotations around axes in the equatorial plane with a site-dependent angle $\theta^{(i)}$ (green shaded area) to imprint the momentum profile of a moving soliton. (c) resp. (d) Charge density profile $\rho(t)/M_{0}$ resp. $\langle\cos{(\beta\hat{\varphi})}\rangle$ as a function of time for $N=101,J=20,\beta^{2}=\pi/20,M_{0}^{\prime}=0.2$, and $v=0$. (e) Collapse of the width $\sigma_{\rho}(t)$ extracted from a Gaussian fit of the charge distribution $\rho$, by rescaling space and time as discussed in the main text. Here we plot $\sigma_{\rho}(t)$ for several $\beta^{2},M_{0}^{\prime}$ and $v=0$. (Inset) Von Neumann entanglement entropy $S_{\text{vN}}$ as a function of time $t$ and the bond that defines the entanglement bipartition for $N=101,J=20,\beta^{2}=\pi/20,M_{0}^{\prime}=0.2$, and $v=0.5$. (f) Collapse of the real part of the vertex operator at half-chain $\langle\cos(\beta\hat{\varphi}(t))\rangle_{\text{mid}}$ for several $\beta^{2},M_{0}^{\prime}$ at $v=0$. The vertical dashed line indicates the “decay” time $t_{\text{d}}$ of the initial soliton, defined by the zero crossing of $\langle\cos(\beta\hat{\varphi}(t))\rangle_{\text{mid}}$ (dash-dotted line).
@@
From here on and in the rest of this work, we perform simulations deep in the gapped phase, such that the values of $\beta^{2},M_{0}^{\prime}$ can be determined from the microscopic parameters $\chi,\lambda_{\kappa}^{\prime},V_{\text{nn}}^{\prime}$, and $\kappa=1$ according to Eq. (9)  [^3] We estimate the maximal relative error in neglecting the $\beta^{2}$-renormalization to $\approx 9\%$.. We also note that the non-commutativity of the two sets of unitaries induces a small systematic effect on the desired profiles in the central region, of size proportional to $(\gamma M_{0}^{\prime})^{-1}$, where the phase profile connects two degenerate minima of the cosine potential [for more details see App. F.1.] We notice that the magnitude of this effect increases with $M_{0}^{\prime}$ and $v$, while decreasing with $\beta^{2}$, which is consistent with the dependence of the momentum profile on the sG parameters. We find that the expectation value of $\hat{J}_{z}$ displays instead a perfect agreement with the theoretical prediction for $\pi_{\text{s}}$.
We emphasize that after applying the gates, in both cases the resulting state displays the same vertex-vertex correlation function (upon accounting for the correct phase of the spin operators $\hat{J}_{\pm}\rightarrow\hat{J}_{\pm}e^{\mp i\beta\varphi_{\text{s}}}$), $\langle\hat{\varphi}^{(i)}\hat{\varphi}^{(j)}\rangle\propto\langle\hat{J}_{+}^{(i)}\hat{J}_{-}^{(j)}\rangle_{\text{c}}=\langle\hat{J}_{+}^{(i)}\hat{J}_{-}^{(j)}\rangle-\langle\hat{J}_{+}^{(i)}\rangle\langle\hat{J}_{-}^{(j)}\rangle$, as the original ground state of the lattice Hamiltonian. In other words, we have effectively changed the one-point functions – allowing us to interpret the resulting state as a semi-classical soliton – while the two-point correlators remain untouched and display the correct vacuum quantum fluctuations. Note that this preparation protocol assumes local programmability of the quantum simulator, similar to Ref. [38]. We further comment on experimentally achievable numerical values of the required parameters in App. F.2.
@@ IV.3 Evolving a semi-classical soliton
We continue by investigating the time evolution of the semi-classical soliton state for both $v=0$ and $v\neq 0$ by evolving the system with the sG Hamiltonian Eq. (8) for $d=1$. Numerically, the time evolution is performed using time-evolving block decimation (TEBD), i.e., applying discrete Trotter steps and updating the MPS representation of the state. We find that TEBD with a moderate maximal bond dimension of $\chi=128$, truncation cutoff $10^{-9}$, and time step $\tau=0.01$ leads to quantitatively converged results [see the discussion in App. F.2 resp. F.3 for more details about static resp. moving soliton simulations]. The following results are obtained from spin lengths $J\in\{16,18,20\}$ for a sufficiently large system size of $N=101$. In the presented parameter regimes our observables are essentially converged in $N$, but require an extrapolation in $J$.
We track the evolution of the initial soliton by monitoring the topological charge density $\rho=\langle\hat{\rho}\rangle$, as well as the expectation value of the interaction potential $1-\langle\cos{(\beta\hat{\varphi})\rangle}$. As time progresses, we find that both observables are increasingly well-approximated by an appropriately normalized Gaussian function characterized by a mean position $\bar{x}$ and a width $\sigma$, measured in lattice units [see App. F], which we discuss next.
@@ IV.3.2 Delocalization: impact of quantum fluctuations
On top of the “classical” motion of the mean soliton position, we find that both the charge density $\rho$ and the cosine term $1-\langle\cos{(\beta\hat{\varphi})}\rangle$ delocalize over time. This is quantified by the width $\sigma(t)$ of a Gaussian fit to both observables. Fig. 5(e) shows the spreading of the width $\sigma_{\rho}(t)$ of the charge density $\rho$ over time. We find the evolution to be self-similar, which is revealed by rescaling the axes $\sigma_{\rho}\rightarrow\sigma_{\rho}\cdot\gamma M_{0}$ and $t\rightarrow t\cdot\beta M_{0}/\gamma^{3/2}$, such that the trajectories of several $\beta^{2},M_{0}^{\prime},$ and $v$ collapse. Here we only present results for moving solitons up to a speed of $v=0.5$. For larger velocities ($v\geq 0.7$), our state preparation creates high-energy excitations at the lattice scale, which propagate with $v\approx c=1$, and interfere with the targeted soliton dynamics.
@@
Here, $v_{\text{spr}}\approx 0.273\cdot\beta$ is the spreading velocity, which we find to exhibit a dominant dependence on $\beta$, with a prefactor estimated from the smallest $M_{0}^{\prime}=0.15$ simulated. The dynamics of $1-\langle\cos{(\beta\hat{\varphi})}\rangle$ shows a similar behaviour with $\sigma_{\text{cos}}(0)=\sigma_{\text{cos},0}\approx 0.794/M_{0}$. Given this evolution, together with the (approximate) conservation of the spatial integral of both quantities in time, the maximum of $1-\langle\cos{(\beta\hat{\varphi})}\rangle$ – which coincides with the value in the middle of the chain – decreases over time [see Fig. 5(f)], and eventually becomes smaller than 1 at some time $t_{\text{d}}$. In other words, after $t_{\text{d}}$ the expectation value $\langle\cos{(\beta\hat{\varphi})}\rangle=\Re{\langle\hat{J}_{+}\rangle/\sqrt{J(J+1)}}$ assumes only positive values in the middle of the chain and the extracted classical phase profile $\langle\beta\hat{\varphi}\rangle=\arg{\langle\hat{J}_{+}\rangle}$ gets restricted to the interval $[-\pi/2,\pi/2]$. This is incompatible with a classical soliton profile $\beta\varphi_{\text{s}}\in[0,2\pi]$ that is expected to smoothly connect the vacua at 00 and $2\pi$. In this sense, the interpretation as a single classical soliton becomes inapplicable and $t_{\text{d}}\approx 4.96/(M_{0}\beta)$ indicates a time where the profile has “decayed” into fluctuations.
@@ IV.3.3 Semi-classical model
[…]
This model thus reproduces the numerically observed spreading of the width upon identifying $\sigma_{v}$ with $v_{\text{spr}}$, and upon assuming that quantum fluctuations are proportional to $\beta$. In App. F.5 we further show that the direct proportionality $\sigma_{v}\propto\beta$ can be attributed to the $zz$-correlations of the quantum ground state. The additional $\sigma_{\delta}$ term in the time-dependent width $\sigma(t)$ can be neglected, as it would simply modify the static profile of the observables, which we instead impose to agree with the classical predictions.
The case $v\neq 0$ can be understood analogously to the static $v=0$ case by interpreting the moving soliton as being static in an inertial frame of reference, which is moving with velocity $v$ with respect to the observer [for further details see App. F.4]. Thus, upon taking into account the relativistic Lorentz contraction of space and dilation of time, we obtain additional Lorentz factors $\gamma$. Moreover, we find that the standard deviation of the velocity for the moving soliton is given by $\sigma_{v}\propto\beta/\sqrt{\gamma}$. Altogether, our semi-classical model also reproduces the $\gamma$-dependence of the rescaled space and time axes in Fig. 5(e), giving further credence to the interpretation of the initial state as a single-soliton wave-packet with the desired velocity.
@@ V Soliton-antisoliton scattering
In the following section, we investigate the real-time dynamics of a counter-propagating soliton-antisoliton pair. After briefly reviewing the expected scattering shifts, we present our numerical results in agreement with the theoretical predictions.
@@ V.1 Quantum and classical scattering shifts
Beyond a single soliton, the classical sine-Gordon model admits several other exact solutions, including the antisoliton $\varphi_{\bar{\text{s}}}$, with topological charge $Q=-1$, as well as a scattering soliton-antisoliton pair, with topological charge $Q=0$. The latter is given by
@@
In the quantum sG model, asymptotic scattering of a soliton and antisoliton with rapidities ±$\pm\theta$ leads to a phase shift $\phi(\theta)=-i\log{S(\theta)}$ determined by the scattering matrix $S(\theta)$ [39]. In the limit $\beta^{2}\rightarrow 0$, the scattering becomes completely transmissive and we restrict our attention to the transmissive part $S_{T}(\theta)$ [69]
@@ V.2 Numerical results
with $\bar{x}_{\text{L}}$ resp. $\bar{x}_{\text{R}}$ the (anti-)soliton position in the left resp. right half of the chain. We find that this approach captures well the (anti-)soliton positions as localized quasi-particles for the essentially free evolution before and after the scattering event. Conversely, we can identify the scattering region in space-time as a single region of high energy density when soliton and antisoliton cannot be distinguished [see Fig. 6(a)]. We thus fit the trajectories in the regions of linear motion and extrapolate the position shift $\delta x$ as the distance between soliton and antisoliton at the scattering time, i.e., the time at which the linear fits cross [see App. G for more details, including a systematic error analysis of this method].
@@
Figure 6: Soliton-antisoliton scattering. (a) Heatmap of the excess lattice energy density $\bar{\varepsilon}_{i}$ as a function of lattice site $i$ and time $t$, with quasi-particle trajectories indicated by the black solid lines. To estimate the scattering position shift, we fit the trajectories with a linear function in the pre-/post-collision time intervals delimited by the vertical dashed black lines. The resulting fits are represented by the dashed red line (soliton) and the blue line (antisoliton). The simulation parameters are $N=101,J=20,\beta^{2}=\pi/13,M_{0}^{\prime}=0.15,v=0.5,\Delta x_{\text{s}\bar{\text{s}}}/a=40$. (b) Scattering position shift $\delta x/a$ as a function of $M_{0}^{\prime}$ for $\beta^{2}=\pi/13$. The circles resp. diamonds are the numerical results for an initial separation $\Delta x_{\text{s}\bar{\text{s}}}/a=40$ resp. $\Delta x_{\text{s}\bar{\text{s}}}/a=50$ sites. The dashed color lines represent the classical position shift according to Eq. (29). The black crosses resp. plus represent the theoretical predictions of Eq. (32) for $\Delta x_{\text{s}\bar{\text{s}}}/a=40$ resp. $\Delta x_{\text{s}\bar{\text{s}}}/a=50$, using as input the renormalized $\beta^{2},M_{0}^{\prime}$ values from the equilibrium analysis and the effective velocity extracted from the trajectories after the scattering event.
@@
The resulting value of the position shift $\delta x$ is shown in Fig. 6(b), which exhibits only slight variations (up to $\mathcal{O}(1)$ lattice site) between different initial soliton-antisoliton separations $\Delta_{\text{s}\bar{\text{s}}}/a=40,50$. We attribute these deviations to the difficulty in precisely locating the quasi-particles. When comparing our numerical results to the theoretical quantum prediction $\delta x_{\text{theo}}=\delta x(\theta,-\theta)$ [Eq. (32)], we use the renormalized values of $\beta^{2}$ and $M_{0}^{\prime}$ determined from the equilibrium analysis in Sec. III and the quasi-particle velocities extracted from the trajectories after the scattering event. Overall, we find good agreement of the numerically obtained position shifts $\delta x$ with the theoretical predictions.

### v2

@@ IV.2 Preparing a semi-classical soliton
To achieve the desired initial state, we first calculate the ground state of the sG Hamiltonian $\hat{H}_{\text{sG}}$ at finite $\beta^{2}$. For our microscopic implementation (at $\kappa=1$), this state corresponds to a correlated chain of spins pointing in +x+x-direction, as $\langle\beta\hat{\varphi}\rangle=\arg{(\langle\hat{J}_{+}\rangle/\sqrt{J(J+1)})}=0$. For a static soliton with $v=0$ and $\gamma=1$, it is now sufficient to imprint the desired phase profile by applying site-dependent $z$-rotations $\hat{U}_{\text{p}}^{(i)}=\exp{\left(-i\beta\varphi_{\text{s}}^{(i)}\hat{J}_{z}^{(i)}\right)}$, with $\varphi_{\text{s}}^{(i)}=\varphi_{\text{s}}(x=i\cdot a,t=0)$ the value of the phase profile at site $i$ [see Fig. 5(a)].
Figure 5: Soliton preparation and dynamics. (a)-(b) Preparation of the semi-classical soliton state: (a) we first apply site-dependent $z$-rotations with an angle $\beta\varphi_{\mathrm{s}}^{(i)}$ (green shaded area) to rotate in the equatorial plane the spins (initially pointing in +x+x-direction). (b) We then apply rotations around axes in the equatorial plane with a site-dependent angle $\theta^{(i)}$ (green shaded area) to imprint the momentum profile of a moving soliton. (c) resp. (d) Charge density profile $\rho(t)/M_{0}$ resp. $\langle\cos{(\beta\hat{\varphi})}\rangle$ as a function of time for $N=101,J=20,\beta^{2}=\pi/20,M_{0}^{\prime}=0.2$, and $v=0$. (e) Collapse of the width $\sigma_{\rho}(t)$ extracted from a Gaussian fit of the charge distribution $\rho$, by rescaling space and time as discussed in the main text. Here we plot $\sigma_{\rho}(t)$ for several $\beta^{2},M_{0}^{\prime}$ and $v=0$. (Inset) Von Neumann entanglement entropy $S_{\text{vN}}$ as a function of time $t$ and the bond that defines the entanglement bipartition for $N=101,J=20,\beta^{2}=\pi/20,M_{0}^{\prime}=0.2$, and $v=0.5$. (f) Collapse of the real part of the vertex operator at half-chain $\langle\cos(\beta\hat{\varphi}(t))\rangle_{\text{mid}}$ for several $\beta^{2},M_{0}^{\prime}$ at $v=0$. The vertical dashed line indicates the “decay” time $t_{\text{d}}$ of the initial soliton, defined by the zero crossing of $\langle\cos(\beta\hat{\varphi}(t))\rangle_{\text{mid}}$ (dash-dotted line).
@@
From here on and in the rest of this work, we perform simulations deep in the gapped phase, such that the values of $\beta^{2},M_{0}^{\prime}$ can be determined from the microscopic parameters $\chi,\lambda_{\kappa}^{\prime},V_{\text{nn}}^{\prime}$, and $\kappa=1$ according to Eq. (9)  [^3] We estimate the maximal relative error in neglecting the $\beta^{2}$-renormalization to $\approx 9\%$.. We also note that the non-commutativity of the two sets of unitaries induces a small systematic effect on the desired profiles in the central region, of size proportional to $(\gamma M_{0}^{\prime})^{-1}$, where the phase profile connects two degenerate minima of the cosine potential [for more details see App. G.1.] We notice that the magnitude of this effect increases with $M_{0}^{\prime}$ and $v$, while decreasing with $\beta^{2}$, which is consistent with the dependence of the momentum profile on the sG parameters. We find that the expectation value of $\hat{J}_{z}$ displays instead a perfect agreement with the theoretical prediction for $\pi_{\text{s}}$.
We emphasize that after applying the gates, in both cases the resulting state displays the same vertex-vertex correlation function (upon accounting for the correct phase of the spin operators $\hat{J}_{\pm}\rightarrow\hat{J}_{\pm}e^{\mp i\beta\varphi_{\text{s}}}$), $\langle\hat{\varphi}^{(i)}\hat{\varphi}^{(j)}\rangle\propto\langle\hat{J}_{+}^{(i)}\hat{J}_{-}^{(j)}\rangle_{\text{c}}=\langle\hat{J}_{+}^{(i)}\hat{J}_{-}^{(j)}\rangle-\langle\hat{J}_{+}^{(i)}\rangle\langle\hat{J}_{-}^{(j)}\rangle$, as the original ground state of the lattice Hamiltonian. In other words, we have effectively changed the one-point functions – allowing us to interpret the resulting state as a semi-classical soliton – while the two-point correlators remain untouched and display the correct vacuum quantum fluctuations. Note that this preparation protocol assumes local programmability of the quantum simulator, similar to Ref. [55]. We further comment on experimentally achievable numerical values of the required parameters in App. G.2.
@@ IV.3 Evolving a semi-classical soliton
We continue by investigating the time evolution of the semi-classical soliton state for both $v=0$ and $v\neq 0$ by evolving the system with the sG Hamiltonian Eq. (8) for $d=1$. Numerically, the time evolution is performed using time-evolving block decimation (TEBD), i.e., applying discrete Trotter steps and updating the MPS representation of the state. We find that TEBD with a moderate maximal bond dimension of $\chi=128$, truncation cutoff $10^{-9}$, and time step $\tau=0.01$ leads to quantitatively converged results [see the discussion in App. G.2 resp. G.3 for more details about static resp. moving soliton simulations]. The following results are obtained from spin lengths $J\in\{16,18,20\}$ for a sufficiently large system size of $N=101$. In the presented parameter regimes our observables are essentially converged in $N$, but require an extrapolation in $J$ to retrieve quantitative results.
We track the evolution of the initial soliton by monitoring the topological charge density $\rho=\langle\hat{\rho}\rangle$, as well as the expectation value of the interaction potential $1-\langle\cos{(\beta\hat{\varphi})\rangle}$. As time progresses, we find that both observables are increasingly well-approximated by an appropriately normalized Gaussian function characterized by a mean position $\bar{x}$ and a width $\sigma$, measured in lattice units [see App. G], which we discuss next.
@@ IV.3.2 Delocalization: impact of quantum fluctuations
On top of the “classical” motion of the mean soliton position, we find that both the charge density $\rho$ and the cosine term $1-\langle\cos{(\beta\hat{\varphi})}\rangle$ delocalize over time. This is quantified by the width $\sigma(t)$ of a Gaussian fit to both observables. Fig. 5(e) shows the spreading of the width $\sigma_{\rho}(t)$ of the charge density $\rho$ over time. We find the evolution to be self-similar, which is revealed by rescaling the axes $\sigma_{\rho}\rightarrow\sigma_{\rho}\cdot\gamma M_{0}$ and $t\rightarrow t\cdot\beta M_{0}/\gamma^{3/2}$, such that the trajectories of several $\beta^{2},M_{0}^{\prime},$ and $v$ collapse. Here we only present results for moving solitons up to a speed of $v=0.5$. For larger velocities ($v\geq 0.7$), our state preparation creates high-energy excitations at the lattice scale, which propagate with $v\approx c=1$, and interfere with the targeted soliton dynamics. We note that a small admixtures of these excitations is visible for all velocities, but their effect is negligible for small $v$.
@@
Here, $v_{\text{spr}}\approx 0.273\cdot\beta$ is the spreading velocity, which we find to exhibit a dominant dependence on $\beta$, with a prefactor estimated from the smallest $M_{0}^{\prime}=0.15$ simulated. The dynamics of $1-\langle\cos{(\beta\hat{\varphi})}\rangle$ shows a similar behaviour with $\sigma_{\text{cos}}(0)=\sigma_{\text{cos},0}\approx 0.794/M_{0}$. Given this evolution, together with the (approximate) conservation of the spatial integral of both quantities in time, the maximum of $1-\langle\cos{(\beta\hat{\varphi})}\rangle$ – which coincides with the value in the middle of the chain – decreases over time [see Fig. 5(f)], and eventually becomes smaller than 1 at some time $t_{\text{d}}$. In other words, after $t_{\text{d}}$ the expectation value $\langle\cos{(\beta\hat{\varphi})}\rangle=\Re{\langle\hat{J}_{+}\rangle/\sqrt{J(J+1)}}$ assumes only positive values in the middle of the chain and the extracted classical phase profile $\langle\beta\hat{\varphi}\rangle=\arg{\langle\hat{J}_{+}\rangle}$ gets restricted to the interval $\pi/2,\pi/2]$. This is incompatible with a classical soliton profile $\beta\varphi_{\text{s}}\in[0,2\pi]$ that is expected to smoothly connect the vacua at 0 and $2\pi$. In this sense, the interpretation as a single classical soliton becomes inapplicable and $t_{\text{d}}\approx 4.96/(M_{0}\beta)$ indicates a time where the profile has “decayed” into fluctuations.
@@ IV.3.3 Semi-classical model
We choose the distribution $W$ as two independent normalized Gaussian distributions in $\delta$ and $v$, centered at $\mu_{\delta}\leavevmode\nobreak\ =\leavevmode\nobreak\ \mu_{v}\leavevmode\nobreak\ =\leavevmode\nobreak\ 0$ and with widths $\sigma_{\delta},\sigma_{v}$, which account for the initial quantum fluctuations. To obtain analytical results, we further approximate both observables $\rho_{\text{s}}(x,t)$ and $1-\cos{(\beta\varphi_{\text{s}}(x,t))}$ at $t=0$ with a Gaussian distribution of width $\sigma_{0}$. Assuming $v\leavevmode\nobreak\ \ll\leavevmode\nobreak\ 1,\gamma\leavevmode\nobreak\ \approx\leavevmode\nobreak\ 1$, these choices lead to evolved observables that remain well-approximated by a Gaussian distribution with a time-dependent width $\sigma(t)=\sqrt{(\sigma_{0}^{2}+\sigma_{\delta}^{2})+(\sigma_{v}\cdot t)^{2}}$ [for more details see App. G].
This model thus reproduces the numerically observed spreading of the width upon identifying $\sigma_{v}$ with $v_{\text{spr}}$, and upon assuming that quantum fluctuations are proportional to $\beta$. In App. G.5 we further show that the direct proportionality $\sigma_{v}\propto\beta$ can be attributed to the $zz$-correlations of the quantum ground state. The additional $\sigma_{\delta}$ term in the time-dependent width $\sigma(t)$ can be neglected, as it would simply modify the static profile of the observables, which we instead impose to agree with the classical predictions.
The case $v\neq 0$ can be understood analogously to the static $v=0$ case by interpreting the moving soliton as being static in an inertial frame of reference, which is moving with velocity $v$ with respect to the observer [for further details see App. G.4]. Thus, upon taking into account the relativistic Lorentz contraction of space and dilation of time, we obtain additional Lorentz factors $\gamma$. Moreover, we find that the standard deviation of the velocity for the moving soliton is given by $\sigma_{v}\propto\beta/\sqrt{\gamma}$. Altogether, our semi-classical model also reproduces the $\gamma$-dependence of the rescaled space and time axes in Fig. 5(e), giving further credence to the interpretation of the initial state as a single-soliton wave-packet with the desired velocity.
@@ V Soliton-antisoliton scattering
In the following section, we investigate the real-time dynamics of a counter-propagating soliton-antisoliton pair. After briefly reviewing the expected scattering shifts, we present our numerical results in quantitative agreement with the theoretical predictions.
@@ V.1 Quantum and classical scattering shifts
Beyond a single soliton, the classical sine-Gordon model admits several other exact solutions, including the antisoliton $\varphi_{\bar{\text{s}}}$, with topological charge Q=1Q=-1, as well as a scattering soliton-antisoliton pair, with topological charge $Q=0$. The latter is given by
@@
In the quantum sG model, asymptotic scattering of a soliton and antisoliton with rapidities $\pm\theta$ leads to a phase shift $\phi(\theta)=-i\log{S(\theta)}$ determined by the scattering matrix $S(\theta)$ [38]. In the limit $\beta^{2}\rightarrow 0$, the scattering becomes completely transmissive and we restrict our attention to the transmissive part $S_{T}(\theta)$ [85]
@@ V.2 Numerical results
with $\bar{x}_{\text{L}}$ resp. $\bar{x}_{\text{R}}$ the (anti-)soliton position in the left resp. right half of the chain. We find that this approach captures well the (anti-)soliton positions as localized quasi-particles for the essentially free evolution before and after the scattering event. Conversely, we can identify the scattering region in space-time as a single region of high energy density when soliton and antisoliton cannot be distinguished [see Fig. 6(a)]. We thus fit the trajectories in the regions of linear motion and extrapolate the position shift $\delta x$ as the distance between soliton and antisoliton at the scattering time, i.e., the time at which the linear fits cross [see App. H for more details, including a systematic error analysis of this method].
@@
Figure 6: Soliton-antisoliton scattering. (a) Heatmap of the excess lattice energy density $\bar{\varepsilon}_{i}$ as a function of lattice site $i$ and time $t$, with quasi-particle trajectories indicated by the black solid lines. To estimate the scattering position shift, we fit the trajectories with a linear function in the pre-/post-collision time intervals delimited by the vertical dashed black lines. The resulting fits are represented by the dashed red line (soliton) and the blue line (antisoliton). The simulation parameters are $N=101,J=20,\beta^{2}=\pi/13,M_{0}^{\prime}=0.15,v=0.5,\Delta x_{\text{s}\bar{\text{s}}}/a=40$. (b) Scattering position shift $\delta x/a$ as a function of $M_{0}^{\prime}$ for $\beta^{2}=\pi/13$. The circles resp. diamonds are the numerical results for an initial separation $\Delta x_{\text{s}\bar{\text{s}}}/a=40$ resp. $\Delta x_{\text{s}\bar{\text{s}}}/a=50$ sites. The dashed color lines represent the classical position shift according to Eq. (29). The black crosses resp. plus represent the theoretical predictions of Eq. (32) for $\Delta x_{\text{s}\bar{\text{s}}}/a=40$ resp. $\Delta x_{\text{s}\bar{\text{s}}}/a=50$, using as input the renormalized $\beta^{2},M_{0}^{\prime}$ values from the equilibrium analysis and the effective velocity extracted from the trajectories after the scattering event. The numerical results are extrapolated to the continuum limit with the procedure described in Sec. II.2.
@@
The resulting value of the position shift $\delta x$ extrapolated to the continuum limit is shown in Fig. 6(b), which exhibits only slight variations (up to $\mathcal{O}(1)$ lattice site) between different initial soliton-antisoliton separations $\Delta_{\text{s}\bar{\text{s}}}/a=40,50$. We attribute these deviations to the difficulty in precisely locating the quasi-particles. When comparing our numerical results to the theoretical quantum prediction $\delta x_{\text{theo}}=\delta x(\theta,-\theta)$ [Eq. (32)], we use the renormalized values of $\beta^{2}$ and $M_{0}^{\prime}$ determined from the equilibrium analysis in Sec. III and the quasi-particle velocities extracted from the trajectories after the scattering event. Overall, we find good agreement of the numerically obtained position shifts $\delta x$ with the theoretical predictions.